OpenStudy (dtan5457):

If f(x)=2x-5 and g(x)=(1+3x)/(5-2x) compute g(f^-1(z))

OpenStudy (freckles):

well first solve y=2x-5 for x

OpenStudy (dtan5457):

y=(x+5)/2

OpenStudy (freckles):

cool \[f^{-1}(x)=\frac{x+5}{2} \\ \text{ so } f^{-1}(z)=\frac{z+5}{2}\]

OpenStudy (freckles):

\[g(\frac{z+5}{2})=?\]

OpenStudy (freckles):

just replace all of those x's in that g thingy with (z+5)/2

OpenStudy (dtan5457):

i got 17+3z/(-2z)

OpenStudy (freckles):

\[\frac{1+3\frac{z+5}{2}}{5-2 \frac{z+5}{2}} \\ \frac{1+\frac{3(z+5)}{2}}{5-(z+5)} \\ \frac{1+\frac{3(z+5)}{2}}{5-z-5}\] did you get that far?

OpenStudy (freckles):

\[\frac{1+\frac{3(z+5)}{2}}{-z}\] yep what you have seems right good job

OpenStudy (dtan5457):

thanks again